Bob wants to make 750 mL of a .250 M sodium carbonate solution. What volume of a 3.50 M stock solution of sodium carbonte should he dilue to 750 mL? How many moles of Na+ are in the 750 mL of diluted solution?
Answer #1

M1V1 = M2V2

3.5 V1 = 0.25 * 750

V1 = (0.25*750)/3.5 = 53.57 mL

Volume of 3.5 M stock solution of sodium carbonte = 53.57 mL

No fo moles of Na+ in solution = 0.25*0.75 = 0.1875 moles